Idkname:

An unopened cereal box contains `15/28` cubic foot of cereal. If the box is full, and the length and height of the box are `5/7` and `1 and (1)/(2)` feet, respectively, what is the width of the box? A. `1/2` foot B. `1/7` foot C. `3/5` foot D. `4/7` foot

Vocaloid:

@dude

dude:

The volume of a box is just length \(\times\) width \(\times\) height They gave you all but 1 value, so you can substitute the missing value into the equation \(V=l\times w\times h\) (First convert 1\(\frac{1}{2}\) into an improper fraction first) \(\dfrac{15}{28}=\dfrac{5}{7}\times \dfrac{3}{2}\times w\) You can solve from width from here